Reports 1-1 of 1 Clear search Modify search
VIS (BS)
fabian.arellano - 12:20 Wednesday 30 June 2021 (17338) Print this report
BS heigth measurement

With Hirata-san.

See pictures in the album BS Remedying work after O3.

Preparation

We released the IP, F0 and set the three keystones to nominal positions.

  • The current position of the F0 keystone and temperature  are -1800 um and T = 23.6 °C respectively (O3GK T = 23 °C).
  • According to klog 16856, we adjusted the ballast mass to have the F0 keystone at 1250 um when the temperature was 22.9 °C. Now that that the temperature is high the keystone moved down a lot.
  • The lifted the F0 keystone with the fishing rod but it reached its upper limit triggering the limit switch. We had to use coil magnet actuation to set it in zero.
  • We set the BF and F1 to zero with coil-magnet actuation.

Measurement

  • The RM does not have scratch markings as the SR RMs.
  • At the HR side of the optic, the RM has vented screws at the same height as the centre of the optic and we can use them as reference.
  • At the AR side of the optic the screws are not at the same height as the centre of the optic, so we cannot use them as reference.
  • With the exception of one marking, the reference height markings made by Terada-san and Tomaru-san on the outer frame and pilars are suitable for inspecting the RM from the AR side, where the screws are not at the samer height as the centre of the optic. We used that one marking to check the height of the optic.
  • We were unable to use rulers becuase the centre of the optic is too high and we don't have long rulers right now.
  • We aligned the laser level and took pictures of it on the RM.
  • The diameter of the vent hole in the screw is 2 mmm and we can use it as a length reference to estimate the height from the picture.
    • At the right side screw (HR,-X,-Y) the RM is lower than the reference by 3 mm approximately.
    • At the left side screw (HR,+X,+Y) the RM is lower than the reference by slightly more then 3 mm.

When the GAS filter keystones are at zero, the BS optic is approximatelly 3 mm below the nominal position.

Images attached to this report
Comments to this report:
fabian.arellano - 16:22 Wednesday 27 October 2021 (18710) Print this report

With Hirata-san.

Summary: the BS optic is 2.8 mm below the average height of the flanges when F0 and BF are close to zero and F1 is alledgely close to zero.

See pictures in the album BS Remedying work after O3.

  • Temperature: T_ondotori = 23.6  °C, T_PT100 = 23.6  °C.
  • We set the F0 anbd BF GAS fiters at zero with coil magnet actuation:
    • F0: 14,000 cnt,
    • BF: 4000 cnt,
  • F1 LVDT is not working well (entry 18191), so we checked the amount of actuation we used the last time we measured the height, when the LVDT was working, and used the same value (entry 17338 from the 30th of June, 2021). This is fine because we have the same temperature today (23.6 °C) and the fishing rod has not been used:
    • F1 actuation: -10,000 cnt (per Data Viewer).
    • F1 LVDT: 787 um (nominal is zero but we have a faulty LVDT primary coil).
  • As a reference we used Tomaru-san's marking of the average of the heights of the four flanges of the chamber.
  • The measured height to be approximately 2.8 mm below the reference.
fabian.arellano - 13:19 Thursday 28 October 2021 (18720) Print this report

With Hirata-san, Terrence and Yano-san.

Summary: We raised the pre-isolator by approximately 2.8 mm and then measured the optic to be 0.3 mm below the height reference.  At the time of the measurement F0 and BF were close to zero and F1 was alledgely close to zero.

See pictures in the album BS Remedying work after O3.

  • Temperature: T_ondotori = 23.5 °C, T_PT100 = 23.7 °C.
  • According to entry  18710, the mirror was 2.8 mm below the reference, therefore, we raised the jacks (M42 X 2mm thread) by rotating them counterclockwise 1.4 turns approximately. The reference is the average height of the centres of the four flanges.
  • Then we set the F0 and BF GAS fiters close to zero and F1 close to 787 um with coil magnet actuation:
    • F0: 13,000 cnt,
    • F1: 10,000 cnt (LVDT readout 788 um becuase there's a problem with the LVDT, real nominal is zero),
    • BF: 4000 cnt.
  • Then we measured the height of the optic: it was 0.3 mm below the reference, which is close enough given the methods we are using for the measurement.
  • We proceeded with leveling the IP by slightly adjusting the jacks. Bringing L close to its setpoint moved L away from its sentpoint, so we settled with having T closer to the setpoint and do further adjustment with the horizontal fishing rods.
    • L: -710 um,
    • T: 100 um.
Search Help
×

Warning

×